A 1200 kg station wagon is moving along a straight highway at 120 m/s. Another car, with

mass 1800 kg and speed 20 m/s, has its centre of mass 40 m ahead of the centre of mass of

the station wagon. Find the position of the centre of mass of the system consisting of the

two automobiles. Find the magnitude of the total momentum of the system. Find the speed

of the system using the speed of the centre of mass.


1
Expert's answer
2022-02-16T15:05:13-0500

"x=\frac{lm_2}{m_1+m_2}=24~m,"

"p=\frac{m_1v_1+m_2v_2}2=90~\frac{t\cdot m}s,"

"v=\frac{v_1+v_2}2=70~\frac ms."